A plane travels576km with the wind in 1.5hours return against the wind in 2 hours. Find the speed of the plane and the speed of the wind?
Answer by richwmiller(17219)

r*t=d
with the wind
(r+c)*1.5=576
r*1.5+c*1.5=576
against the wind
r*2-c*2=576
1) r*2-c*2=576
2) r*1.5+c*1.5=576
multiply 1) by 3 multiply 2) by 4
r*6.0+c*6.0=2304
r*6-c*6=1728
add
r*12.0=4032
r=336.0
r=336 mph plane
336*1.5+c*1.5=576
504.0+c*1.5=576
c*1.5=72.0
c=72.0/1.5
c=48.0 mph wind
check
(336+c)*dt=576
(336+48)*1.5=576
(384)*1.5=576
576.0=576
(336-c)*ut=576
(336-48)*2=576
(288)*2=576
576=576
ok
or
more simply
r*t=d
with the wind
(r+c)*1.5=576
r+c=576/1.5
r+c=384.0
against the wind
r-c=576/2
r-c=288
we have
r-c=288
r+c=384.0
add
2r=672.0
r=336.0
r=336 mph plane
336*1.5+c*1.5=576
504.0+c*1.5=576
c*1.5=72.0
c=72.0/1.5
c=48.0 mph wind
check
(336+c)*dt=576
(336+48)*1.5=576
(384)*1.5=576
576.0=576
(336-c)*ut=576
(336-48)*2=576
(288)*2=576
576=576
ok